OpenStudy (anonymous):

if a and b are the measures of two first quadrant angles and sin a = 4/5 and sin b = 5/13, find sin a+b

OpenStudy (dumbcow):

angle addition property \[\sin(a+b) = \sin a \cos b + \sin b \cos a\]

OpenStudy (dumbcow):

also \[\sin^{2} \theta + \cos^{2} \theta = 1\] this is how you can find cos(a) and cos(b)

OpenStudy (anonymous):

I tried to to that but got (20/39)+(1/3) or 11/13. The available answers are (63/65) (33/65) (16/65) (56/65)

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

looks like you didn't get the right values for cos(a) and cos(b) cos a = 3/5 cos b = 12/13

OpenStudy (anonymous):

So it would be sin(a+b)=(4/5)(12/13)+(3/5)(5/13) = 63/65?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Thanks!

OpenStudy (dumbcow):

try to go back and see what you did wrong in finding the cosine values